OpenStudy (akgorgeous):

2/3 (x − 2) = 4x Solve for x.

OpenStudy (jackthegreatest):

2/3 x - 4/3 = 4x multiply both sides by 3 2x - 4 = 12x -4 = 10 x finish it up @akgorgeous

OpenStudy (studygurl14):